OpenStudy (anonymous):

Use the inverse matrix method to solve: x+2y -z=10 -2x +3y+z=6 3x-2y+2z=-19 Select one: a. |1 -3 5| b. |-1 3 -5| c. |1 3 5| d. no solution

OpenStudy (kirbykirby):

If you let \[ \left[ \begin{array}{ccc} 1 & 2 & -1 \\ -2 & 3 & 1 \\ 3 & -2 & 2 \end{array} \right]\] be the matrix \(A\), then you can write your system of equations as: \( \left[ \begin{array}{ccc} 1 & 2 & -1 \\ -2 & 3 & 1 \\ 3 & -2 & 2 \end{array} \right]\) \( \left[ \begin{array}{c} x\\y\\z \end{array} \right]\) =\( \left[ \begin{array}{c} 10\\6\\-19 \end{array} \right]\) or \(A\vec{x}=\vec{b}\) So multiplying by the inverse \(A^{-1}\): \(A^{-1}A\vec{x}=A^{-1}\vec{b}\\ \vec{x}=A^{-1}\vec{b} \) Now this requires you to invert the matrix \(A\). Can you find the inverse?

OpenStudy (anonymous):

is it d?

OpenStudy (kirbykirby):

no

OpenStudy (anonymous):

uhhh b?

OpenStudy (kirbykirby):

are you just guessing letters? I just hope you tried attempting it

OpenStudy (anonymous):

i am im doing it now

OpenStudy (kirbykirby):

:)

OpenStudy (anonymous):

A?

OpenStudy (kirbykirby):

what did you find as the inverse for A?

OpenStudy (anonymous):

2

OpenStudy (kirbykirby):

the inverse of a matrix is again a matrix :o

OpenStudy (anonymous):

ugg this is hard im really bad at math

OpenStudy (kirbykirby):

explaining how to invert a matrix on here will be very long to type :S

OpenStudy (anonymous):

the best i can go with is C?

OpenStudy (anonymous):

its looks correct

OpenStudy (kirbykirby):

I don't get C

OpenStudy (anonymous):

im going with A

OpenStudy (kirbykirby):

I feel like you are just guessing letters :(

OpenStudy (anonymous):

im not im just frasturated

OpenStudy (kirbykirby):

The link there shows you all the steps on how to invert a matrix. The last stepis to multiply your inverted matrix with the vector \(\vec{b}\) using matrix multiplication

OpenStudy (anonymous):

i got it..........

OpenStudy (kirbykirby):

oh great